7. Two fair cubical dice are thrown: one is red and one is blue. The random variable M represents the score on the red die minus the score on the blue die. (a) Find the distribution of M. (b) Write down E(M). (c) Find Var(M).


1
Expert's answer
2021-06-22T08:57:11-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Sample & Red,Blue & M \\
\hline
 1 & 1,1 & 0 \\
 \hdashline
 2 & 1,2 & -1 \\
 \hdashline
 3 & 1,3 & -2 \\
 \hdashline
 4 & 1,4 & -3\\
 \hdashline
 5 & 1,5 & -4 \\
\hdashline
 6 & 1,6 & -5 \\
 \hdashline
7 & 2,1 & 1 \\
 \hdashline
8 & 2,2 & 0 \\
 \hdashline
9 & 2,3 & -1 \\
 \hdashline
10 & 2,4 & -2\\
 \hdashline
 11 & 2,5 & -3 \\
\hdashline
 12 & 2,6 & -4 \\
\hdashline
13 & 3,1 & 2 \\
 \hdashline
14 & 3,2 & 1 \\
 \hdashline
15 & 3,3 & 0 \\
 \hdashline
16 & 3,4 & -1\\
 \hdashline
17 & 3,5 & -2 \\
\hdashline
18 & 3,6 & -3 \\
 \hdashline
19 & 4,1 & 3 \\
 \hdashline
20 & 4,2 & 2 \\
 \hdashline
 21 & 4,3 & 1 \\
 \hdashline
22 & 4,4 & 0\\
 \hdashline
23 & 4,5 & -1 \\
\hdashline
24 & 4,6 & -2 \\
 \hdashline
25 & 5,1 & 4 \\
 \hdashline
26 & 5,2 & 2 \\
 \hdashline
 27 & 5,3 & 2 \\
 \hdashline
 28 & 5,4 & 1\\
 \hdashline
 29 & 5,5 & 0 \\
\hdashline
 30 & 5,6 & -1 \\
 \hdashline
31 & 6,1 & 5 \\
 \hdashline
 32 & 6,2 & 4 \\
 \hdashline
 33 & 6,3 & 3 \\
 \hdashline
 34 & 6,4 & 2\\
 \hdashline
 35 & 6,5 & 1 \\
\hdashline
 36 & 6,6 & 0 \\
 
 \hline
\end{array}"




"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 M & f & f(M) & Mf(M)& M^2f(M) \\ \hline
 -5 & 1 & 1/36 & -5/36 & 25/36 \\
 \hdashline
 -4 & 2 & 2/36 & -8/36 & 32/36 \\
 \hdashline
 -3 & 3 & 3/36 & -9/36 & 27/36 \\
 \hdashline
 -2 & 4 & 4/36 & -8/36 & 16/36 \\
 \hdashline
 -1 & 5 & 5/36 & -5/36 & 5/36 \\
 \hdashline
 0 & 6 & 6/36 & 0 & 0 \\
 \hdashline
 1 & 5 & 5/36 & 5/36 & 5/36 \\
 \hdashline
 2 & 4 & 4/36 & 8/36 & 16/36 \\
 \hdashline
 3 & 3 & 3/36 & 9/36 & 27/36 \\
 \hdashline
 4 & 2 & 2/36 & 8/36 & 32/36 \\
 \hdashline
 5 & 1 & 1/36 & 5/36 & 25/36 \\
 \hdashline
 Total & 36 & 1 & 0 & 35/6 \\ \hline
\end{array}"



1.


"\begin{matrix}
 M & p(M) \\
\hline 
 -5 & 1/36 \\
-4 & 2/36 \\
 -3 & 3/36 \\
-2 & 4/36 \\
 -1 & 5/36 \\
0 & 6/36 \\
 1 & 5/36 \\
2 & 4/36 \\
 3 & 3/36 \\
4 & 2/36 \\
 5 & 1/36 \\
\end{matrix}"


2.


"E(M)=\sum Mp(M)=0"

The mean of the sampling distribution of the sample means is equal to the

the mean of the population.


"E(\bar{X})=0=\mu"



3.


"Var(M)=\sum M^2p(M)-(\sum Mp(M))^2"




"=\dfrac{35}{6}-(0)^2=\dfrac{35}{6}"